An art dealer bought 12 etchings for $375. He sold some of them at $35 each and the rest at $40 each, making a profit of $80 on the transaction. How many etchings did he sell at each price?
Answer by richwmiller(17219)

x+y=12,
35x+40y=375+80
x = 5 at 35 , y = 7 at 40